### v1

# Correspondence between two exact gravitational lens equations in a static and spherically symmetric spacetime

Abstract
Virbhadra and Ellis have proposed an exact equation (referred to as VE equation) for the gravitational lens in a static and spherically symmetric spacetime [Phys. Rev. D 62,084003 (2000)], whereas an improved equation (referred to as BTOA equation) has been derived by Bozza [Phys. Rev. D 78, 103005 (2008)] and later by Takizawa, Ono and Asada [Phys. Rev. D, 102, 064060 (2020)]. VE and BTOA equations seem to be very different from each other. The present paper shows that there exists an unphysical branch in VE equation. Consequently, VE equation can be reduced by removing the unphysical branch. The reduced version of VE equation is found to be the same as BTOA equation when a suitable transformation is made between the deflection angles defined differently in the two formulations. An explicit expression of the transformation is found. We also argue possible numerical errors when the transformation between the deflection angles is ignored.
pacs:
04.40.-b, 95.30.Sf, 98.62.Sb
I Introduction
The gravitational lens plays crucial roles in modern cosmology and gravitational physics SEF ; Petters ; Dodelson ; Keeton ; Will , where most of earlier works focus on weak deflection of light. However, the direct imaging of the massive black hole candidates by the Event Horizon Telescope has drastically changed the situation EHT2019 ; EHT2022 . Exact (or very accurate) treatments are needed not only for theoretical studies Darwin ; Bozza ; Perlick but also for observational researches on the gravitational lens in the strong gravity region. Notably, the project of the next generation Event Horizon Telescope is increasing the importance of the lens by strong gravity ngEHT .
As a pioneering work, Virbhadra and Ellis (VE) have proposed an exact equation for the gravitational lens in a static and spherically symmetric spacetime VE . Their geometrical approach is almost exact, though they assume that the deflection of light is caused on the lens plane. Their assumption is a good approximation when the observer, lens and source configuration is almost symmetric. Recently, VE equation has been used for detailed discussions of image distortions of Schwarzschild lensing Virbhadra2022 ; Virbhadra2024 .
Instead of assuming such a symmetrical configuration, Bozza has made use of the Ohanian relation for a more general configuration Ohanian to obtain an improved version of the exact lens equation Bozza2008 . Takizawa, Ono and Asada have later rediscovered this equation by using a state-of-art formulation GW ; Ishihara2016 ; Ishihara2017 based on the Gauss-Bonnet theorem Takizawa2020 . For its simplicity, the improved equation is referred to as BTOA equation. BTOA equation seems to be very different from VE equation. Is it true?
The main purpose of this paper is to examine whether VE equation is really different from BTOA one. First, we shall show that VE equation has a duality that it has not only a physical branch but also an unphysical one. Hence, VE equation can be reduced by removing the unphysical branch. Secondly, we shall argue that the reduced version of VE equation can be the same as BTOA equation when we make a suitable transformation between the deflection angles defined differently in the two formulations.
This paper is organized as follows. Section II discusses the unphysical branch in VE equation. Section III clarifies a relation between the reduced VE equation and BTOA one. Section IV argues possible numerical errors when the deflection angle transformation is ignored. Throughout this paper, we focus on a static and spherically symmetric spacetime.
II Two exact lens equations
II.1 VE and BTOA equations
Let us begin with briefly summarizing VE and BTOA equations. Both equations describe the mapping between the source angle $\beta$ and image angle $\theta$ in terms of angular diameter distances in rigorous manners (e.g. See Figure 1). $D_{L}$, $D_{S}$ and $D_{LS}$ denote the angular diameter distances between the observer and the lens plane, between the observer and the source plane, and between the lens and the source plane, respectively.
A crucial difference between the two equations is how to treat the deflection angles of light. See Figure 1. VE equation assumes a symmetric configuration in which the observer and source are equidistant from the lens, as usually assumed in the conventional lens equation with the thin lens approximation SEF .
Figure 1: The deflection occurs at the point P on the lens plane. The deflection angle $\alpha_{VE}$ is assumed to be positive in this figure.
In Figure 1, we consider a relation among the lengths $IS$, $SU$ and $IU$ on the source plane as $IS=IU-SU$. From this relation, VE equation can be obtained as
$D_{S}\tan\beta=D_{S}\tan\theta-D_{LS}[\tan\theta+\tan(\alpha_{VE}-\theta)],$ (1)
where $\alpha_{VE}$ is an angle measured at the point P on the lens plane in Figure 1. From the aspect of the triangular inequality, Dabrowski and Schunck pointed out difficulties of using VE equation DS .
In order to avoid an almost symmetric assumption, Ohanian considered the tangent line at the source and that at the observer (See Figure 2) and the intersection point Q Ohanian . Note that the intersection point Q is generally off the lens plane. For the quadrilateral OLSQ, a relation among the four interior angles can be obtained. It is Ohanian relation Ohanian . Bozza rewrote it in terms of the angular diameter distances to arrive at an improved version of an exact lens equation as
$D_{S}\tan\beta=\frac{D_{L}\sin\theta-D_{LS}\sin(\alpha_{B}-\theta)}{\cos(\alpha_{B}-\theta)},$ (2)
where $\alpha_{B}$ is defined at the intersection point Q in Figure 2.
Figure 2: The deflection point Q in the BTOA equation is off the lens plane.
By using the Gauss-Bonnet method on the gravitational lens GW , another exact lens equation is derived as Takizawa2020
$\alpha_{TOA}-\theta-\arcsin\left(\frac{D_{L}}{\sqrt{(D_{LS})^{2}+(D_{S})^{2}\tan^{2}\beta}}\sin\theta\right)$
$+\arctan\left(\frac{D_{S}}{D_{LS}}\tan\beta\right)$
$=0,$ (3)
where $\alpha_{TOA}$ is defined at the intersection point Q in Figure 2. This equation is equivalent to Eq. (2). See Reference Takizawa2020 for a proof of the equivalence. Hence, $\alpha_{B}=\alpha_{TOA}$.
II.2 Duality in VE equation
By straightforward calculations, Eq. (1) is rewritten into a quadratic in $\tan\theta$ as
$D_{L}\tan\alpha_{VE}(\tan\theta)^{2}+D_{S}(1-\tan\beta\tan\alpha_{VE})\tan\theta$
$-(D_{S}\tan\beta+D_{LS}\tan\alpha_{VE})=0.$ (4)
In Figure 1, we may expect that a positive $\tan\theta$ existed uniquely for a given $\alpha_{VE}$. However, it is not the case. Eq. (1) is equivalent to Eq. (4) and it thus admits two $\tan\theta$. One of the two $\tan\theta$ corresponds to Figure 1, whereas the other $\tan\theta$ is unphysical because it does not describe the light ray in Figure 1.
For $\alpha_{VE}$ in Figure 1, Eq. (4) is formally solved for $\tan\theta$ as
$\tan\theta=\frac{D_{S}(\tan\beta\tan\alpha_{VE}-1)\pm\sqrt{\cal D}}{2D_{L}\tan\alpha_{VE}},$ (5)
where
${\cal D}\equiv$ $(D_{S})^{2}(\tan\beta\tan\alpha_{VE}-1)^{2}$
$+4D_{L}\tan\alpha_{VE}(D_{S}\tan\beta+D_{LS}\tan\alpha_{VE}).$ (6)
By direct calculations, ${\cal D}$ can be rewritten as
${\cal D}=A\left(\tan\alpha_{VE}-\frac{B}{A}\right)^{2}+\frac{4D_{L}D_{LS}(D_{S})^{2}}{A\cos^{2}\beta},$ (7)
where $A\equiv(D_{S})^{2}\tan^{2}\beta+4D_{L}D_{LS}$, $B\equiv D_{S}(D_{LS}-D_{L})\tan\beta$, and $A>0$. Eq. (7) tells ${\cal D}>0$ for any $\alpha_{VE}$.
The plus and minus signs in Eq. (5) reflect two branches in VE equation. Note that Eq. (5) does not mean a solution for the lens equation for which $\alpha_{VE}$ is not a given number but it should be treated as a function of $\theta$. Eq. (5) is still an equation for $\theta$ as shown later.
II.3 Unphysical configuration for VE equation
In this subsection, we shall examine what is the unphysical branch in VE equation. First, we consider another configuration of the observer, lens and source system shown by Figure 3, where $\alpha_{VE2}$ denotes the deflection angle. We note that $0<\alpha_{VE}<\pi$ and $-\pi<\alpha_{VE2}<0$ in Figures 1 and 3, respectively. For the strong deflection of a relativistic image, $2N\pi<\alpha_{VE}<(2N+1)\pi$ in Figure 1 and $-(2N+1)\pi<\alpha_{VE2}<-2N\pi$ in Figure 3 for $N=1,2,\cdots$, where $N$ denotes the winding number of the light ray. We consider the deflection of light, for which $\alpha_{VE}\neq 0$ and $\alpha_{VE2}\neq 0$.
Figure 3: Another deflection point $P^{\prime}$ is located on the lens plane. Here, $\alpha_{VE2}$ is negative.
Here, let us set $\alpha_{VE2}=\alpha_{VE}-\pi$ ($\alpha_{VE2}=\alpha_{VE}-(4N+1)\pi$ for a pair of relativistic images with the winding number $N$), which leads to $\tan\alpha_{VE2}=\tan\alpha_{VE}$. This means that the different configuration specified by $\alpha_{VE2}(\neq\alpha_{VE})$ satisfies the same lens equation as Eq. (1). Therefore, Eq. (1) for $\tan\theta$ has two solutions, where one solution corresponds to Figure 1 and the other describes Figure 3. The angle $\theta$ for the physical configuration in Figure 1 is greater than $\theta$ for the unphysical configuration in Figure 3.
In the weak deflection of light, the image in Figure 1 is called the primary image on the same side of the intrinsic source position with respect to the lens SEF . Hence, $\tan\theta$ in Figure 1 is greater than that in Figure 3. Up to this point, we have focused on $\alpha_{VE}>0$.
Next, we consider the case of $\alpha_{VE}<0$, for which Figure 3 corresponds to a physical configuration, whereas Figure 1 is unphysical. In the weak deflection, therefore, the image in Figure 3 is the secondary image on the opposite side of the intrinsic source position with respect to the lens SEF .
II.4 Duality in VE solutions
Next, let us classify two branches in VE equation. For our graphical method, it is convenient to rewrite Eq. (1) into
$\frac{D_{L}}{D_{LS}}\tan\theta-\frac{D_{S}}{D_{LS}}\tan\beta=\frac{\tan\alpha_{VE}-\tan\theta}{1+\tan\alpha_{VE}\tan\theta},$ (8)
where $D_{S}-D_{LS}=D_{L}$ is used.
It is worthwhile to note that $\tan\theta$ satisfying Eq. (8) corresponds to an intersection point between a straight line as
$y=ax-b,$ (9)
and a hyperbolic curve as
$y=\frac{c-x}{1+cx},$ (10)
where $x\equiv\tan\theta$, $a\equiv D_{L}/D_{LS}\geq 0$, $b\equiv D_{S}\tan\beta/D_{LS}$, and $c\equiv\tan\alpha_{VE}$.
In the rest of the paper, we assume $\tan\beta\geq 0$ without loss of generality, because the spacetime under study is spherically symmetric. Hence, $b\geq 0$. On the other hand, the sign of $c$ is arbitrary, because $-\infty<\tan\alpha_{VE}<+\infty$. $c=\tan\alpha_{VE}\neq 0$ because the light is deflected. Hence, $c>0$ or $c<0$.
There are four possible types of intersections:
(1) For $c>0$, see Figure 4 for schematic plots of Eqs. (9) and (10). This figure shows that $x(=\tan\theta)$ of one intersection is positive, whereas the other is negative. The former and latter intersections correspond to Figures 1 and 3, respectively.
Figure 4: $c\equiv\tan\alpha_{VE}>0$. There are two solutions of VE equation . The red (in color) line denotes Eq. (9) and the blue (in color) curves mean Eq. (10). They have two intersection points. One of the intersections is $x>0$ and the other is $x<0$.
(2) For $-b<c<0$, see Figure 5. $x$ for both solutions is positive. Only for this case, there are no solutions corresponding to Figure 3.
Figure 5: $-b<c<0$, where $b=D_{S}(\tan\beta)/D_{LS}$. There are two solutions of VE equation. The red (in color) line denotes Eq. (9) and the blue (in color) curves mean Eq. (10). They have two intersection points. Both of the intersections are $x>0$.
(3) For $-b=c<0$, one solution for $x$ is positive and the other is $x=0$. However, the latter case is corresponding to the lens direction $\theta=0$. However, this direction is not observable and thus can be ignored.
(4) For $c<-b$, see Figure 6. One solution is positive and the other is negative.
Figure 6: $c<-b$. There are two solutions of VE equation. The red (in color) line denotes Eq. (9) and the blue (in color) curves mean Eq. (10). They have two intersection points. One of the intersections is $x>0$ and the other is $x<0$.
In schematic plots in Figures 4-6, $\tan\theta$ at an intersection is a real number.
Note that all of the above (1)-(4) are not necessarily realized. Care is needed for the case (2) as follows. In Figure 1, we may expect that a positive $\tan\theta$ existed uniquely for given $\alpha_{VE}$. However, the number of positive $\tan\theta$ in the case (2) is two. This means that one $\tan\theta$ corresponds to Figure 1, while the other $\tan\theta$ is unphysical because it does not correspond to Figure 1.
In deriving Eq. (1), moreover, Figure 1 is assumed, where $\tan\theta\geq 0$ is implicitly assumed. For this setup, therefore, the negative $x$ in the above cases (1) and (4) must be excluded. Therefore, there can exist an unphysical branch in Eq. (1).
We mention the other case, $\tan\theta<0$. This case cannot be described by Eq. (1) when $\alpha_{VE}$ is the same as that for $\tan\theta\geq 0$. Namely, Figure 1 should be replaced by Figure 3, which leads to another VE equation with another $\alpha_{VE2}$ indicated in Figure 3. $\alpha_{VE2}$ does not equal to $\alpha_{VE}$.
II.5 Reduced VE equation
There are plus and minus signs in Eq. (5). Which sign should be chosen? The answer relies upon the sign of $\alpha_{VE}$. We consider $\alpha_{VE}\neq 0$ because we discuss the nonzero deflection of light.
(1)$\alpha_{VE}>0$:
This case corresponds to Figure 1. The plus sign in Eq. (5) should be thus chosen as
$\tan\theta=\frac{D_{S}(\tan\beta\tan\alpha_{VE}-1)+\sqrt{\cal D}}{2D_{L}\tan\alpha_{VE}}.$ (11)
(2)$\alpha_{VE}<0$:
This case means Figure 3. The minus sign in Eq. (5) should be thus chosen as
$\tan\theta=\frac{D_{S}(\tan\beta\tan\alpha_{VE}-1)-\sqrt{\cal D}}{2D_{L}\tan\alpha_{VE}}.$ (12)
III Correspondence between VE and BTOA equations
III.1 Transformation between the VE and BTOA deflection angles
For the same configuration of the lens, observer and source with given $\beta$ and $\theta$, the reduced VE and BTOA equations should equal to each other. For Eq. (11) to equal to Eq. (2), a relation between $\alpha_{VE}$ and $\alpha_{B}$ must be
$\tan\alpha_{VE}=\frac{D_{S}F(E-F\tan\beta)}{D_{LS}F^{2}+D_{S}EF\tan\beta-D_{L}E^{2}},$ (13)
where functions $E$ and $F$ of $\alpha_{B}$ are defined as
$E$ $\equiv D_{LS}\sin\alpha_{B}+D_{S}\tan\beta\cos\alpha_{B},$ (14)
$F$ $\equiv D_{L}+D_{LS}\cos\alpha_{B}-D_{S}\tan\beta\sin\alpha_{B}.$ (15)
Figure 7 shows a relation between $\alpha_{VE}$ and $\alpha_{B}$ in Eq. (13).
It is worthwhile to mention that BTOA equation as Eq. (2) can be simply rewritten in terms of $E$ and $F$ as
$\tan\theta=\frac{E}{F}.$ (16)
Figure 7: $\tan\alpha_{VE}$ vs $\tan\alpha_{B}$. Top: $\beta=0$, and $D_{LS}/D_{L}=$ 1 (blue), 0.5 (orange), 0.3 (green) and 0.1 (red). Bottom: The equidistance case. $\beta=$ 0 (blue), 0.1 (orange), 0.2 (green) and 0.3 (red).
IV Numerical errors by ignoring the deflection angle transformation
First, we consider two cases $D_{L}\ll D_{LS}$ and $D_{L}\gg D_{LS}$. For $D_{L}\ll D_{LS}$, Eq. (13) is expanded in terms of $D_{L}/D_{LS}$ as
$\tan\alpha_{VE}-\tan\alpha_{B}=O\left(\frac{D_{L}}{D_{LS}}\right).$ (17)
Hence, $\alpha_{VE}$ is very close to $\alpha_{B}$. This means that the substitution of $\alpha_{B}$ into $\alpha_{VE}$ in Eq. (1) can be a good approximation even if the transformation is neglected.
For $D_{L}\gg D_{LS}$, on the other hand, Eq. (13) is expanded in terms of $D_{LS}/D_{L}$ as
$\tan\alpha_{VE}-\tan\alpha_{B}=-\frac{1}{\tan\beta\cos\alpha_{B}}+O\left(\frac{D_{LS}}{D_{L}}\right).$ (18)
This implies that the difference between $\alpha_{VE}$ and $\alpha_{B}$ can be significantly large in the limit as $D_{LS}/D_{L}\to 0$, especially for nearly coaligned configurations $\beta\sim 0$. This case of $D_{L}\gg D_{LS}$ and $\beta\sim 0$ are corresponding to the observation of photons from galactic centers, notably a vicinity of a photon surface of the massive black hole candidate in Sgr A.
Figure 8: $\tan\theta_{VE}$ and $\tan\theta_{B}$. Here, the transformation of $\alpha_{VE}$ is ignored when we solve the reduced VE equation. The dashed line corresponds to $\tan\theta_{VE}=\tan\theta_{B}$ as a reference. $D_{LS}/D_{L}=10^{-1}$ (blue), $10^{-2}$ (red), and $10^{-3}$ (green), where $\beta=0.1$ is chosen.
How large numerical errors are caused when we ignore the transformation between the two deflection angles? For this purpose, we intend to equate $\alpha_{VE}$ and $\alpha_{B}$. We focus on the case of $|\tan\alpha_{VE}|\ll 1$ and $|\tan\alpha_{B}|\ll 1$. From Eqs. (1) and (2), then, we can obtain
$\cfrac{\tan\theta_{VE}-\tan\beta}{\cfrac{D_{LS}}{D_{S}}(1+\tan^{2}\beta)}=\cfrac{\tan\theta_{B}-\tan\beta}{\cfrac{D_{LS}}{D_{S}}+\tan^{2}\beta}+O(\tan^{2}\alpha_{VE},\tan^{2}\alpha_{B}),$ (19)
where $\theta_{VE}$ and $\theta_{B}$ denote a solution for Eqs. (1) and (2), respectively. We ignore the second order in $\tan\alpha_{B}$ below.
For $\tan\theta_{VE}$ to equal to $\tan\theta_{B}$, the two denominators in Eq. (19) must be the same as each other. However, it is not the case. See Figure 8 for numerical plots of Eq. (19).
The two denominators are very close to each other, if and only if $|\beta|\ll 1$ and $D_{L}\ll D_{LS}$. This condition means that the observer, lens and source are almost coaligned and the observer is close to the lens. This condition does not meet in most astronomical situations.
The exceptional case is the gravitational lens by the Sun. In many astronomical cases ($\beta\neq 0$ and $D_{L}\gg D_{LS}$), conversely speaking, the ignorance of the transformation between the deflection angles would result in a significant difference between $\tan\theta_{VE}$ and $\tan\theta_{B}$.
See Figure 9 for numerical calculations of the difference when the transformation is ignored. The numerical plots in this figure are consistent with an expression as
$\tan\alpha_{VE}-\tan\alpha_{B}=\frac{D_{L}}{D_{LS}}\sin^{2}\beta\tan\alpha_{B}+O\left(\tan^{2}\alpha_{B}\right),$ (20)
which can be obtained from Eq. (13) in the approximation as $|\tan\alpha_{B}|\ll 1$. Eq. (20) is consistent with also Figure 7 in the domain of small $\tan\alpha_{B}$.
Figure 9: Numerical errors by ignoring the transformation between $\alpha_{VE}$ and $\alpha_{B}$ when the reduced VE equation is solved. The parameters correspond to the Sgr A Gillessen in the approximation of the Schwarzschild lens, where the angular radius of the black hole shadow PT is about 26 microarcsec., $D_{L}=8$ kpc, and $D_{LS}/D_{L}=1.2\times 10^{-10}$ EHT2022 . Here, $\delta\equiv|\theta_{VE}/\theta_{B}-1|$ is the relative difference due to neglecting the transformation.
V Summary
We have shown that there exists an unphysical branch in VE equation. VE equation is reduced by removing the unphysical branch. The reduced VE equation has been found to be the same as BTOA equation when the transformation between the deflection angles is suitably done. We have argued possible numerical errors when the transformation is ignored. In particular, numerical errors can be significant in the use of the VE equation for the Sgr A.
It is left for future to extend the current result to a stationary and axisymmetric lens.
Acknowledgements.
We are grateful to Keita Takizawa and Marcus Werner for fruitful discussions. We thank Yuuiti Sendouda and Ryuichi Takahashi for useful conversations. This work was supported in part by Japan Science and Technology Agency (JST) SPRING, Grant Number, JPMJSP2152 (R.K.) and in part by Japan Society for the Promotion of Science (JSPS) Grant-in-Aid for Scientific Research, No. 20K03963 and 24K07009(H.A.).